# A perspective on inelastic light scattering spectroscopy for probing transport of collective acoustic excitations

### v1

@@ IV  Applications of ISS
Figure 4: (a) Measured (solid circles) and calculated (line) dispersion relation of the lowest frequency surface acoustic mode in a silicon thin film (Inset: dispersion curves of the four lowest frequency modes) Maznev and Every (2009a). Reprinted from A. A. Maznev et al. Appl. Phys. Lett. 95, 011903 (2009), with the permission of AIP Publishing. (b) TGS signal from a micro-patterned 1D phononic structure, showing acoustic oscillations associated with non-leaky, long-lived surface acoustic modes. (Inset: the same signal over a longer time scale) Maznev and Wright (2009). Reprinted from A. A. Maznev et al. J. Appl. Phys. 105, 123530 (2009), with the permission of AIP Publishing. (c) Acoustic dispersion of a 2D granular crystal, where red and black markers represent the measured frequencies with and without silica microspheres, respectively. A horizontal dotted line indicates the resonance frequency arising from microsphere-substrate adhesion Boechler et al. (2013). Reprinted figure with permission from N. Boechler et al. Interaction of a Contact Resonance of Microspheres with Surface Acoustic Waves, Phys. Rev. Lett. 111, 036103 (2013). Copyright 2013 by the American Physical Society. (d) Measured spin helix lifetimes versus wavevector in GaAs/AlGaAs quantum wells, showing enhanced lifetime peaks to higher wavevectors with increasing doping asymmetry Koralek et al. (2009). Reproduced with permission from J. D. Koralek et al. Emergence of the persistent spin helix in semiconductor quantum wells, Nature 458, 7238 (2009). Copyright 2009 Springer Nature.
@@
In ISS, surface acoustic waves (SAWs) are launched by a density grating from electrostriction or a thermal grating from optical absorption, depending on the nature of the medium. The wavelength of the excited SAW is determined by the optical grating period, enabling non-contact measurements of SAWs and guided modes in thin films, multilayers, liquids, and interfaces Maznev and Every (2009b); Kim et al. (2017); Maznev and Wright (2009); Simmonds et al. (2024); Yang and Nelson (1995); Yan et al. (1988); Paolucci and Nelson (2000). By measuring dispersion and velocities of surface acoustic modes, elastic moduli, film thickness, and depth profiling of structural defects can be obtained Rogers et al. (2000); Steigerwald et al. (2009); Mechri et al. (2009). Figure 4(a) shows experimental results on stacked crystalline Si with polycrystalline Cu-Ta thin film layers, where the lowest acoustic mode was observed to exhibit a negative group velocity at a frequency near 190 MHz, analogous to Lamb modes in free plates Maznev and Every (2009a). ISS can also provide access to interfacial acoustics; for instance, co-propagating surface modes and a short-lived, high-velocity interfacial mode were observed at TiO2-water interfaces Kasinski et al. (1989), whereas at solid and glass-forming liquid interfaces, it has been applied to capture structural relaxation dynamics during the glass transition Glorieux et al. (2006). Another application of ISS includes probing the quality of adhesion and mechanical properties of coatings, such as MnO2 films, where SAW velocity dispersion and Scholte wave velocities were measured to estimate Young’s modulus and porosity of coatings Sermeus et al. (2014).
@@
ISS has also been employed to measure attenuation and dispersion relations of SAWs in phononic systems. In Fig. 4(b), a micro-patterned SiO2-Si structure exhibits TGS signal from spatially localized, non-leaky surface acoustic modes with lifetimes of hundreds of nanoseconds Maznev and Wright (2009). On the other hand, grooved silica substrate showed additional band gaps and leaky SAW modes Malfanti et al. (2011). In locally resonant acoustic metamaterials, hybridized flexural Lamb waves Khanolkar et al. (2015) and avoided crossings between contact resonances and Rayleigh modes, shown in Fig. 4(c), were observed Boechler et al. (2013); Eliason et al. (2016); Vega-Flick et al. (2017). More recently, ISS has moved from point spectroscopy to Brillouin microscopy of micromechanics in transparent and soft materials Krug et al. (2022); Li et al. (2024); Krug et al. (2019); Li et al. (2022). Recent work has demonstrated rapid elastography with sub-millisecond acquisition time, enabling 2D Brillouin shift mapping of PDMS–methanol mixture with reduced photodamage Li et al. (2024).
@@
By adjusting pump beams to be cross-polarized, transient spin grating spectroscopy can be realized, which is used to probe spin dynamics of electrons and excitons, and valley polarization dynamics across various semiconductors and magnetic materials, including low-dimensional systems such as quantum wells and quantum dots Wang et al. (2013); Walser et al. (2012); Yang et al. (2012a); Carter et al. (2006); Yang et al. (2012b); Mahmood et al. (2018); Ishiguro et al. (2007); Watanuki et al. (2005); Scholes et al. (2006); Kim et al. (2006); Huxter et al. (2010); Crisp et al. (2013). Early seminal work using transient spin grating demonstrated that electron drift mobility and diffusion in GaAs/AlGaAs multiple quantum wells can be measured via exploiting 90° polarization-rotation of a linearly polarized probe beam diffracted from spin grating, which isolates electron spin transport from ambipolar charge transport Cameron et al. (1996). Wavevector–resolved spin grating subsequently revealed non-diffusive spin-polarization transport in semiconductor quantum wells, with the maximum spin decay time at a finite wavevector that contradicts predictions from simple diffusion theory due to spin-orbit coupling Weber et al. (2007). In addition, spin Coulomb drag effect in a two-dimensional electron gas was demonstrated using spin gratings, where electron–electron scattering transfers momentum between up- and down-spin channels, thereby suppressing macroscopic spin transport Weber et al. (2005). As shown in Fig. 4(d), spin gratings have been used to probe the tunability of spin polarization lifetimes; by varying the level of doping asymmetry to manipulate the strength of Rashba spin-orbit interaction, it was demonstrated that the peak lifetimes of persistent spin helix modes in GaAs/AlGaAs quantum wells can be progressively shifted along wavevector Koralek et al. (2009). More recently, TGS has been extensively used to probe valley polarization dynamics in low-dimensional materials lacking inversion symmetry Mahmood et al. (2018); Wang et al. (2019); Kuhn et al. (2020), including the observation of fluence-dependent valley decay rate in monolayer MoSe2, where strong exciton-exciton interactions lead to fast valley depolarization lifetimes of a few picoseconds Mahmood et al. (2018). Together, these studies highlight the functionality of TGS for probing spin and valley transport across systems from bulk crystals to emerging quantum materials, providing insights into spin-orbit coupling and valley excitons that are essential for developing spintronic, valleytronic, and optoelectronic devices.
@@
At the onset of ballistic phonon transport, the thermal conductivity contribution from heat-carrying, thermal phonons with MFP ($\Lambda$) longer than the thermal length starts to decrease Maznev et al. (2011). The corresponding grating period dependent thermal conductivity is expressed as Minnich (2012)
@@
where $x=q\Lambda_{\omega}$, and K(x)=dS/dxK(x)=-dS/dx with $S(x)$ being the heat suppression function that accounts for effective heat flux, which is reduced compared to Fourier law predicted heat flux due to quasi-ballistic transport Minnich (2012). The cumulative thermal conductivity, $F(\Lambda_{\omega})(=\int_{0}^{\Lambda_{\omega}}f(\Lambda)d\Lambda)$, provides insight into microscopic heat conduction properties by enabling to examine the contributions of MFPs of heat carriers to the thermal conductivity. By changing the integration variable from MFP to frequency Yang and Dames (2013), one can also express Eq. 5 as
@@
where $\mathbf{p}$ indexes the polarization and $C_{\mathbf{p}}$ and $v_{\mathbf{p}}$ are the specific heat and group velocity of thermal energy carriers. In Eq. 6, $\Lambda_{\mathbf{p}}(\omega)$ is MFPs versus frequency, representing scattering mechanisms that depend on the frequency. In TGS, it has been established that the cumulative thermal conductivity and the frequency versus MFPs can be reconstructed by solving Eqs. 5 and 6, and that ill-posed problems of Eqs. 5 and 6 can be solved using numerical optimizations or statistical methods Minnich (2012); Ravichandran et al. (2018); Robbins et al. (2019).
@@
The measurements of the grating period thermal conductivity in macroscopic thin films have provided insights into the understanding of microscopic transport properties governing heat conduction along the in-plane direction. While it has been recognized that atomic scale imperfections such as surface roughness strongly affects phonon boundary scattering, hence the effective thermal resistance, experimentally resolving the boundary conditions at which phonon specular reflections occur has been challenging Chen (2021). Ravichandran et al. have demonstrated a metrology to extract specular parameters from the ISTS measurements on crystalline silicon thin films, as shown in Fig. 5(c) Ravichandran et al. (2018). The experimentally resolved specularity is shown in Fig. 5(d), in which the spectral dependence of phonon reflections at the boundary can be seen.
@@
Additionally, ISTS measurements on thermally conductive polymers Robbins et al. (2019); Kim et al. (2022) have revealed microscopic transport properties that lead to exceptional thermal conductivities. It has been shown that the thermal conductivities of structurally-engineered semi-crystalline polymers including polyethylene (PE) can achieve orders of magnitude increases, which can be as high as those in many condensed matters. While such thermal conductivity enhancements have been typically interpreted using effective medium models Takayanagi et al. (1964); Choy and Young (1977); Lu et al. (2018), debate has existed concerning the origin of the high thermal conductivity, particularly for highly oriented polymers Ronca et al. (2017); Xu et al. (2019). Recent ISTS measurements applied to stretched PEs have provided insights into microscopic origins that lead to high thermal conductivities in these specimens. As shown in Fig. 5(e), for partially oriented PEs, it has been shown that MFP of phonons that substantially contribute to the heat conduction is on the order of 100 nm, which exceeds the grain size of $\sim 60$ nm, indicating ballistic propagation of phonons across the grain boundaries Robbins et al. (2019). Additionally, more recent measurements on highly oriented PEs have shown that thermal conductivity of macroscopic thin films of PE can be as high as $\sim 50$ Wm-1K-1, which was primarily attributed to thermal acoustic phonons that propagate over a few hundreds nanometers, limited by enlarged crystals in them Kim et al. (2022).
@@
ISTS experiments have also been applied to amorphous solids and experimentally probed the scattering of thermal vibrations that occurs distinctly in them, which was not captured using prior numerical studies Kim et al. (2021). An experimental approach that enables to reconstruct the frequency dependent MFPs along with MFP accumulation function by utilizing constraints were demonstrated and applied to amorphous silicon (aSi). Results are shown in Fig 5(f), in which power-law dependence of MFPs ($\Lambda\sim\omega^{-n}$) can be seen, indicating the characteristics of scattering that depends on the frequency. In particular, Rayleigh-type scattering as a primary scattering mechanism ($n=4$) was observed above 1 THz for aSi, which contradicts with predictions from prior normal mode analysis yet agrees with experimental results in other glassy substances such as vitreous silica.
@@ V  Outlook and Conclusion
To make further developments in Brillouin scattering, recent efforts have focused on pushing the limits of the spatiotemporal resolution of the technique. The introduction of short wavelength light sources in TGS setups has enabled grating periods approaching sub-micrometer length scales, allowing researchers to probe non-diffusive dynamics in materials. Free electron lasers in the extreme ultraviolet (EUV) Bencivenga et al. (2015, 2019); Miedaner et al. (2024) and X-ray Rouxel et al. (2021) bands have been implemented to further increase penetration depth and push grating period limits below nanometer scales. These grating periods can elucidate magnetic dynamics and unconventional heat transport at nanometer scales. In the time domain, fast electronic processes such as electron-hole recombination can be temporally resolved with attosecond-scale temporal resolution by implementing both sub-femtosecond excitation and detection processes Fidler et al. (2019); Quintero-Bermudez et al. (2025), providing valuable insights into fast electronic processes such as carrier thermalization and recombination processes. Furthermore, when combined with enhanced temporal and spatial resolution, spin gratings also have potential for studying electronic spin transport dynamics that can establish the theoretical background for spintronic device applications.
Brillouin scattering has also expanded into applications beyond materials science, including optical communications and biological imaging. In photonics, stimulated Brillouin processes have been investigated to exploit enhanced Brillouin gain in optical fibers and waveguides for applications in microwave photonics and signal processing Kobyakov et al. (2010); Wolff et al. (2015); Botter et al. (2022). This enables ultra-low-loss on-chip platforms that provide narrowband optical gain, demonstrating the feasibility of integrated Brillouin photonics Neijts et al. (2024); Ye et al. (2025). In the field of life sciences, Brillouin microscopy has emerged as one of the instrumental techniques for mapping mechanical properties of soft and liquid biological samples. A pioneering work reported the first in situ three-dimensional imaging of the intraocular lens of a mouse using confocal Brillouin microscopy Scarcelli and Yun (2008). Since then, Brillouin technique has been applied to diverse biological systems, including 3D imaging of human cornea both in vivo Scarcelli and Yun (2012) and ex vivo Scarcelli et al. (2013), probing intracellular stress granule mechanics during liquid-to-solid phase transition Antonacci et al. (2018), and mapping intracellular biomechanics within a whole living organism Scarcelli et al. (2015). These advances highlight the possible translation of Brillouin microscopy from basic research to clinical diagnostics, particularly in ophthalmology Shao et al. (2019); Kabakova et al. (2024); Bouvet et al. (2025).

### v2

@@ IV  Applications of ISS
In impulsive stimulated scattering processes, surface acoustic waves (SAWs) are launched by a density grating from electrostriction or a thermal grating from optical absorption, depending on the nature of the medium. The wavelength of the excited SAW is determined by the optical grating period, enabling non-contact measurements of SAWs and guided modes in thin films, multilayers, liquids, and interfaces maznev_jap_2009; kim_aplmat_2017; maznev_japplphys_2009; simmonds_aipadv_2024; yang_jchemphys_1995; yan_jchemphys_1988; paolucci_jchemphys_2000. By measuring dispersion and velocities of surface acoustic modes, elastic moduli, film thickness, and depth profiling of structural defects can be obtained rogers2000optical; steigerwald_apl_2009; mechri_apl_2009. Figure 4(a) shows experimental results on stacked crystalline Si with polycrystalline Cu-Ta thin film layers, where the lowest acoustic mode was observed to exhibit a negative group velocity at a frequency near 190 MHz, analogous to Lamb modes in free plates maznev_apl_2009. ISS can also provide access to interfacial acoustics; for instance, co-propagating surface modes and a short-lived, high-velocity interfacial mode were observed at TiO2-water interfaces kasinski_jcp_1989, whereas at solid and glass-forming liquid interfaces, it has been applied to capture structural relaxation dynamics during the glass transition glorieux_jap_2006. Another application of ISS includes probing the quality of adhesion and mechanical properties of coatings, such as MnO2 films, where SAW velocity dispersion and Scholte wave velocities were measured to estimate Young’s modulus and porosity of coatings sermeus_jap_2014.
@@
ISS has also been employed to measure attenuation and dispersion relations of SAWs in phononic systems. In Fig. 4(b), a micro-patterned SiO2-Si structure exhibits TGS signal from spatially localized, non-leaky surface acoustic modes with lifetimes of hundreds of nanoseconds maznev_japplphys_2009. On the other hand, grooved silica substrate showed additional band gaps and leaky SAW modes malfanti_jmechphyssol_2011. In locally resonant acoustic metamaterials, hybridized flexural Lamb waves khanolkar_apl_2015 and avoided crossings between contact resonances and Rayleigh modes, shown in Fig. 4(c), were observed boechler2013interaction; eliason_apl_2016; vega-flick_prb_2017. More recently, ISS has moved from point spectroscopy to Brillouin microscopy of micromechanics in transparent and soft materials krug_optexp_2022; li_photonres_2024; krug2019impulsive; Jiarui_2022_Optics_ISBSapp1. Recent work has demonstrated rapid elastography with sub-millisecond acquisition time, enabling 2D Brillouin shift mapping of PDMS–methanol mixture with reduced photodamage li_photonres_2024.
@@
By adjusting pump beams to be cross-polarized, transient spin grating spectroscopy can be realized, which is used to probe spin dynamics of electrons and excitons, and valley polarization dynamics across various semiconductors and magnetic materials, including low-dimensional systems such as quantum wells and quantum dots wang_natcommun_2013; walser_natphys_2012; Yang_natphys_2012; carter_prl_2006; yang_prl_2012; Mahmood_2018_NanoLetters_ISBSapp2; ishiguro_apl_2007; watanuki_apl_2005; scholes_prb_2006; kim_jphyschemb_2006; huxter_chemphyslett_2010; crisp_nanolett_2013. Early seminal work using transient spin grating demonstrated that electron drift mobility and diffusion in GaAs/AlGaAs multiple quantum wells can be measured via exploiting 90° polarization-rotation of a linearly polarized probe beam diffracted from spin grating, which isolates electron spin transport from ambipolar charge transport cameron_prl_1996. Wavevector–resolved spin grating subsequently revealed non-diffusive spin-polarization transport in semiconductor quantum wells, with the maximum spin decay time at a finite wavevector that contradicts predictions from simple diffusion theory due to spin-orbit coupling weber_prl_2007. In addition, spin Coulomb drag effect in a two-dimensional electron gas was demonstrated using spin gratings, where electron–electron scattering transfers momentum between up- and down-spin channels, thereby suppressing macroscopic spin transport weber_nature_2005. As shown in Fig. 4(d), spin gratings have been used to probe the tunability of spin polarization lifetimes; by varying the level of doping asymmetry to manipulate the strength of Rashba spin-orbit interaction, it was demonstrated that the peak lifetimes of persistent spin helix modes in GaAs/AlGaAs quantum wells can be progressively shifted along wavevector koralek2009emergence. More recently, TGS has been extensively used to probe valley polarization dynamics in low-dimensional materials lacking inversion symmetry Mahmood_2018_NanoLetters_ISBSapp2; wang_apl_2019; kuhn_laserphot_2020, including the observation of fluence-dependent valley decay rate in monolayer MoSe2, where strong exciton-exciton interactions lead to fast valley depolarization lifetimes of a few picoseconds Mahmood_2018_NanoLetters_ISBSapp2. Together, these studies highlight the functionality of TGS for probing spin and valley transport across systems from bulk crystals to emerging quantum materials, providing insights into spin-orbit coupling and valley excitons that are essential for developing spintronic, valleytronic, and optoelectronic devices.
@@
At the onset of ballistic phonon transport, the thermal conductivity contribution from heat-carrying, thermal phonons with MFP ($\Lambda$) longer than the thermal length starts to decrease Maznev_PRB_2011_supp. The corresponding grating period dependent thermal conductivity is expressed as Minnich_PRL_2012
@@
where $x=q\Lambda_{\omega}$, and K(x)=dS/dxK(x)=-dS/dx with $S(x)$ being the heat suppression function that accounts for effective heat flux, which is reduced compared to Fourier law predicted heat flux due to quasi-ballistic transport Minnich_PRL_2012. The cumulative thermal conductivity, $F(\Lambda_{\omega})(=\int_{0}^{\Lambda_{\omega}}f(\Lambda)d\Lambda)$, provides insight into microscopic heat conduction properties by enabling to examine the contributions of MFPs of heat carriers to the thermal conductivity. By changing the integration variable from MFP to frequency Yang_PRB_2013, one can also express Eq. 5 as
@@
where $\mathbf{p}$ indexes the polarization and $C_{\mathbf{p}}$ and $v_{\mathbf{p}}$ are the specific heat and group velocity of thermal energy carriers. In Eq. 6, $\Lambda_{\mathbf{p}}(\omega)$ is MFPs versus frequency, representing scattering mechanisms that depend on the frequency. In TGS, it has been established that the cumulative thermal conductivity and the frequency versus MFPs can be reconstructed by solving Eqs. 5 and 6, and that ill-posed problems of Eqs. 5 and 6 can be solved using numerical optimizations or statistical methods Minnich_PRL_2012; Navaneetha_2018_PRX_specularity; Andrew_2019_PNAS_DrawRatio.
The measurements of the grating period dependent thermal conductivity in macroscopic thin films have provided insights into the understanding of microscopic transport properties governing heat conduction along the in-plane direction. It should be noted that heat suppression functions have been obtained by solving the phonon Boltzmann transport equation (BTE) within the framework of relaxation time approximation (RTA) or Callaway model to describe lattice thermal conductivity Shiga_PRB_2012; Skelton_PRB_2014; Minnich_PRB_2015. However, these approximations show limitations in describing phonon transport, especially in high thermal conductivity materials and in the non-diffusive heat flow regime where momentum-conserving phonon scattering processes dominate Lindsay_PRB_2013; Jinlong_PRB_2014; Malviya_PRB_2023; Malviya_J.Phys.Condens.Matter_2025. Therefore, symmetry properties of the linearized scattering operator and low-rank representations of the solution have been suggested to accurately solve the BTE with reduced computational cost when the RTA and Callaway assumptions are no longer valid Chaput_PRL_2013; Hua_PRB_2020; Navaneetha_arXiv_2025.
@@
While it has been recognized that atomic scale imperfections such as surface roughness strongly affects phonon boundary scattering, hence the effective thermal resistance, experimentally resolving the boundary conditions at which phonon specular reflections occur has been challenging Chen_NatRevPhys_2021. In estimating specularity parameter, numerical models such as the Ziman’s specularity model Ziman_2001 have failed to fully explain the specularity parameter and often yield contradictory results Cuffe_PRL_2013; Hertzberg_NanoLett_2014; Jiang_PRB_2018 . Furthermore, prior studies on thermal transport in silicon nanowires have reported that the unphysical specularities, which are incompatible with Ziman’s theory, yield the best agreement with the measured temperature-dependent thermal conductivity trend Mingo_PRB_2003; Kazan_JAP_2010. Above problems can be attributed to the fact that while the measured signal represents an ensemble average over the spot size, the Ziman’s model depends on atomic-scale local properties such as surface roughness and impurities, which complicates the analysis. Overall, these imply that it is necessary to experimentally resolve the parameter as a function of phonon wavelength. Ravichandran et al. have demonstrated a metrology to extract specular parameters from the ISTS measurements on crystalline silicon thin films, as shown in Fig. 5(c) Navaneetha_2018_PRX_specularity. The experimentally resolved specularity is shown in Fig. 5(d), in which the spectral dependence of phonon reflections at the boundary can be seen.
@@
Additionally, ISTS measurements on thermally conductive polymers Andrew_2019_PNAS_DrawRatio; Kim_Natcomm_2022 have revealed microscopic transport properties that lead to exceptional thermal conductivities. It has been shown that the thermal conductivities of structurally-engineered semi-crystalline polymers including polyethylene (PE) can achieve orders of magnitude increases, which can be as high as those in many condensed matters. While such thermal conductivity enhancements have been typically interpreted using effective medium models takayanagi_JPSC_1964; choy_polymer_1977; Lu_JAP_2018, debate has existed concerning the origin of the high thermal conductivity, particularly for highly oriented polymers ronca_polymer_2017; Xu_Natcomm_2019. Recent ISTS measurements applied to stretched PEs have provided insights into microscopic origins that lead to high thermal conductivities in these specimens. As shown in Fig. 5(e), for partially oriented PEs, it has been shown that MFP of phonons that substantially contribute to the heat conduction is on the order of 100 nm, which exceeds the grain size of $\sim 60$ nm, indicating ballistic propagation of phonons across the grain boundaries Andrew_2019_PNAS_DrawRatio. Additionally, more recent measurements on highly oriented PEs have shown that thermal conductivity of macroscopic thin films of PE can be as high as $\sim 50$ Wm-1K-1, which was primarily attributed to thermal acoustic phonons that propagate over a few hundreds nanometers, limited by enlarged crystals in them Kim_Natcomm_2022.
@@
ISTS experiments have also been applied to amorphous solids and experimentally probed the scattering of thermal vibrations that occurs distinctly in them, which was not captured using prior numerical studies Kim_2021_PRM_aSi_TG. An experimental approach that enables to reconstruct the frequency dependent MFPs along with MFP accumulation function by utilizing constraints were demonstrated and applied to amorphous silicon (aSi). Results are shown in Fig 5(f), in which power-law dependence of MFPs ($\Lambda\sim\omega^{-n}$) can be seen, indicating the characteristics of scattering that depends on the frequency. In particular, Rayleigh-type scattering as a primary scattering mechanism ($n=4$) was observed above 1 THz for aSi, which contradicts with predictions from prior normal mode analysis yet agrees with experimental results in other glassy substances such as vitreous silica.
@@ V  Outlook and Conclusion
To make further developments in Brillouin scattering, recent efforts have focused on pushing the limits of the spatiotemporal resolution of the technique. The introduction of short wavelength light sources in TGS setups has enabled grating periods approaching sub-micrometer length scales, allowing researchers to probe non-diffusive dynamics in materials. Free electron lasers in the extreme ultraviolet (EUV) Bencivenga_2015_nature_EUV_TGS_1; Bencivenga_2019_scienceAdv_EUV_TGS_2; Miedaner_2024_scienceadv_EUV_TGS_3 and X-ray Rouxel_2021_naturePhotonics_XrayTG bands have been implemented to further increase penetration depth and push grating period limits below nanometer scales. These grating periods can elucidate magnetic dynamics and unconventional heat transport at nanometer scales. In the time domain, fast electronic processes such as electron-hole recombination can be temporally resolved with attosecond-scale temporal resolution by implementing both sub-femtosecond excitation and detection processes Fidler_2019_Natcom_atto1; Bermudez_2025_ACS_atto2, providing valuable insights into fast electronic processes such as carrier thermalization and recombination processes. Furthermore, when combined with enhanced temporal and spatial resolution, spin gratings also have potential for studying electronic spin transport dynamics that can establish the theoretical background for spintronic device applications.
Brillouin scattering has also expanded into applications beyond materials science, including optical communications and biological imaging. In photonics, stimulated Brillouin processes have been investigated to exploit enhanced Brillouin gain in optical fibers and waveguides for applications in microwave photonics and signal processing Kobyakov_advoptphoton_2010; wolff_pra_2015; botter_sciadv_2022. This enables ultra-low-loss on-chip platforms that provide narrowband optical gain, demonstrating the feasibility of integrated Brillouin photonics neijts_aplphot_2024; ye_aplphot_2025. In the field of life sciences, Brillouin microscopy has emerged as one of the instrumental techniques for mapping mechanical properties of soft and liquid biological samples. A pioneering work reported the first in situ three-dimensional imaging of the intraocular lens of a mouse using confocal Brillouin microscopy Scarcelli_2008_NatPhotonics_ISBSapp4. Since then, Brillouin technique has been applied to diverse biological systems, including 3D imaging of human cornea both in vivo Scarcelli_optexp_2012 and ex vivo scarcelli_invopht_2013, probing intracellular stress granule mechanics during liquid-to-solid phase transition Antonacci_communbio_2018, and mapping intracellular biomechanics within a whole living organism Scarcelli_natmet_2015. These advances highlight the possible translation of Brillouin microscopy from basic research to clinical diagnostics, particularly in ophthalmology shao_scirep_2019; Kabakova_natrev_2024; Bouvet_natphot_2025.